OpenStudy (anonymous):

How do you solve the equation -12x - 5 = -3x + 2

OpenStudy (amistre64):

a = b only if a-b=0

OpenStudy (amistre64):

(-12x - 5) = (-3x + 2) (-12x - 5) - (-3x + 2) = 0 -12x - 5 + 3x - 2 = 0